an aircraft covers a distance of 42km in 2minutes and 30 seconds what is the average speed?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

an aircraft covers a distance of 42km in 2minutes and 30 seconds what is the average speed?

2minutes and 30 seconds = 2 and 1/2 minutes = 5/2 minutes= (5/2 )/60
= 5/120 hrs
speed = d/t = 32/ (5/120)
=768 kmph

Answer by ikleyn(53749)

.
an aircraft covers a distance of 42 km in 2 minutes and 30 seconds. what is the average speed?
~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth are incorrect/inaccurate.
        I came to bring a correct accurate solution.


2 minutes and 30 seconds = 2 and 1/2 minutes = 5/2 minutes = (5/2 )/60 = 5/120 of an hour.

speed = d/t = 42 / (5/120) = 1008 kilometers per hour.         ANSWER

Solved correctly.